# Theory of Code Space: Do Code Agents Understand Software Architecture?

### 28 Feb 2026 v1

@@ Abstract
AI code agents excel at isolated tasks—writing functions, fixing localized bugs—yet struggle with complex, multi-file software engineering that requires understanding how dozens of modules relate to each other. We hypothesize that these failures stem from an inability to construct, maintain, and update coherent architectural beliefs during codebase exploration. We introduce Theory of Code Space (ToCS), a benchmark that evaluates this capability by placing agents in procedurally generated codebases under partial observability, requiring them to build structured belief states over module dependencies, cross-cutting invariants, and design intent. Our framework makes three key design choices: (1) a procedural codebase generator producing medium-complexity Python projects with four typed edge categories reflecting different discovery methods—from syntactic imports to config-driven dynamic wiring—with planted architectural constraints and verified ground truth; (2) a partial observability harness where agents explore under a budget, ensuring that exploration strategy matters; and (3) periodic belief probing via structured JSON, producing a time-series of architectural understanding rather than a single final snapshot. We decompose the Active-Passive Gap from spatial reasoning benchmarks into selection and decision components, and introduce Architectural Constraint Discovery as a code-specific evaluation dimension absent from spatial domains. Preliminary experiments with four rule-based baselines and five frontier LLM agents from three providers validate the benchmark’s discriminative power: methods span a wide performance range (F1 from 0.129 to 0.646), LLM agents discover semantic edge types invisible to all baselines, yet weaker models score below simple heuristics—revealing that belief externalization, faithfully serializing internal understanding into structured JSON, is itself a non-trivial capability and a first-order confounder in any belief-probing benchmark. We release the complete toolkit as open-source software and invite the community to evaluate additional models and contribute codebase patterns.111Code: https://github.com/che-shr-cat/tocs
@@ 1 Introduction
Our contributions are:
1.
The ToCS benchmark framework: the first evaluation for active architectural belief construction, revision, and exploitation in code (§3).
2.
A procedural codebase generator producing medium-complexity projects with four typed edge categories, planted invariants, and verified ground truth (§4).
3.
Pilot experiments with four baselines and five frontier LLM agents from three providers, showing that the strongest agent surpasses all baselines (F1 = 0.646 vs. 0.577) while revealing a wide capability spread: LLM agents discover all four edge types but vary 6 in belief externalization fidelity (§6).
4.
An open benchmark with evaluation harness, scoring pipeline, and community call for expanded model evaluation (§7).
@@ 3.2 Environment and Action Space
The agent operates under a budget (default ). Critically, SEARCH returns only locations, never content—ensuring that architectural understanding requires deliberate OPEN decisions.
@@ 3.3 Cognitive Map Probing
The cognitive map contains: (1) component beliefs with status (observed/inferred/unknown), purpose, exported symbols with typed signatures, and typed dependency edges with per-element confidence; (2) invariant beliefs with structured canonical form (type, src, dst, via, pattern) and evidence pointers; (3) uncertainty tracking—explicit list of unexplored regions. The resulting time-series of maps (one every steps) captures how understanding develops, not just the final state.
@@ 3.4 Evaluation Modes
Scope note: v0.1 evaluates the Construct operation in Active mode only. The passive conditions and APG decomposition are implemented in the harness and reserved for future work with frontier model evaluations.
@@ 4.2 Four Edge Types
calls_api (17%): Runtime function calls between modules. Requires reading function bodies.
@@
data_flows_to (7%): Data dependency where one module’s output feeds another’s input. Requires understanding orchestration logic.
These proportions are not tunable parameters but emerge from the architectural template: a Pipeline with stages naturally generates 3–4 more imports edges because every module requires static imports, while semantic edge types arise only at specific architectural boundaries (adapterstage, orchestratorstage, consecutive stages). This mirrors real-world Pipeline codebases, where import edges dominate. Different architectural patterns (event-driven, microservices) would yield different distributions—a key motivation for extending the generator. Crucially, roughly one-third of edges are invisible to import-following (specific to our Pipeline template; other architectural patterns would yield different proportions), creating a meaningful gap between syntactic analysis and semantic understanding.
@@ 4.3 Invariant Planting
Each codebase contains 15–16 planted constraints across five types: boundary (forbidden dependencies), dataflow (required processing chains), interface (access-only-through-ABC), invariant (naming/structural conventions), and purpose (design rationales). Every constraint has a structured canonical form (type, src, dst, via, pattern) for machine-comparable scoring and at least one evidence source (test file, structural pattern, or documentation).
@@ 5 Metrics
Compare predicted edges from the cognitive map against ground truth. An edge matches if (source, target, type) all agree. We report precision, recall, and F1 separately.
@@
Match agent-discovered constraints against planted constraints via structured form comparison—exact field matching on (type, src, dst, via), not text similarity. This eliminates ambiguity in natural-language descriptions.
@@ 6 Experiments
We evaluate four rule-based exploration strategies and five frontier LLM agents from three providers on three generated codebases (seeds 42, 123, 999) under identical conditions: budget , probe interval . These results are preliminary; we report them to validate the benchmark’s discriminative power and invite community replication.
@@ 6.1 Methods
Gemini 3.1 Pro: Google’s frontier pro model (gemini-3.1-pro-preview).
@@
All models receive identical prompts: a system prompt describing the partial observability environment and available actions, per-turn action prompts showing remaining budget and opened files, and structured JSON probe prompts with schema and example (Appendix A). Temperature is set to 0 except for GPT-5.3-Codex (reasoning model constraint). All probe responses were parsed successfully via the deterministic repair parser (zero format failures across all runs).
@@ 6.2 Results
| Oracle | baseline | 1.000 | 1.000 | 1.000 | — | 0.000 | — |
| Config-Aware | baseline | 0.577 | 0.736 | 0.475 | 0.212 | 0.000 | 13 |
| Random | baseline | 0.538 | 1.000 | 0.368 | 0.142 | 0.000 | 13 |
| BFS-Import | baseline | 0.293 | 1.000 | 0.173 | 0.079 | 0.000 | 13 |
| Claude Sonnet 4.6 | LLM | 0.646 | 0.991 | 0.479 | 0.377 | 0.000 | 12 |
| GPT-5.3-Codex | LLM | 0.564 | 0.768 | 0.460 | 0.243 | 0.000 | 14 |
| Gemini 2.5 Flash | LLM | 0.328 | 0.568 | 0.232 | 0.227 | 0.000 | 13 |
| Gemini 3.1 Pro | LLM | 0.262 | 0.894 | 0.156 | 0.135 | 0.000 | 9 |
| Gemini 3 Flash | LLM | 0.129 | 0.713 | 0.071 | 0.120 | 0.000 | 12 |
@@
| Claude Sonnet 4.6 | 0.55 | 0.15 | 0.00 | 1.00 |
| GPT-5.3-Codex | 0.57 | 0.15 | 0.31 | 0.32 |
| Gemini 2.5 Flash | 0.32 | 0.10 | 0.00 | 0.00 |
| Gemini 3.1 Pro | 0.20 | 0.05 | 0.06 | 0.05 |
| Gemini 3 Flash | 0.09 | 0.08 | 0.00 | 0.00 |
@@ 6.3 Analysis
Claude Sonnet 4.6 surpasses baselines.
Claude Sonnet 4.6 achieves F1 = 0.646, surpassing the best rule-based strategy (Config-Aware, 0.577) by 7 points—the first LLM agent to outperform all baselines on this benchmark. It opens 12 files (vs. 13 for baselines) yet achieves higher recall (0.479 vs. 0.475), with near-perfect precision (0.991: only 1 false positive across all 3 codebases). Critically, it discovers all 19 ground-truth REGISTRY_WIRES edges with zero false positives (Table 3)—matching Config-Aware’s recall but without its 38 false positives (Config-Aware achieves only 33% precision on this edge type). This suggests Claude reads both the config file and registry loading logic, correctly inferring which modules are dynamically wired.
@@
Table 3 reveals the most distinctive finding: LLM agents collectively discover all four edge types, while baselines find at most two (IMPORTS and REGISTRY_WIRES for Config-Aware). All five LLMs found CALLS_API edges—invisible to all rule-based baselines—confirming genuine code comprehension beyond syntactic parsing. GPT-5.3-Codex is the only method to discover DATA_FLOWS_TO edges (5 of 16, 31% recall, with perfect precision), demonstrating understanding of how data flows through the pipeline orchestrator. These edges require multi-hop reasoning: tracing one module’s return value through the runner into the next module’s input.
GPT-5.3-Codex: strong but variable.
GPT-5.3-Codex achieves F1 = 0.564 on average (second-highest), but with high variance across codebases: 0.719 on seed 42 vs. 0.392 on seed 999. It discovers the broadest range of edge types (all four), including 5 DATA_FLOWS_TO edges with perfect precision. However, it over-generates REGISTRY_WIRES edges (6 correct, 29 false positives across 3 codebases), reducing its overall precision to 0.768. The high variance may stem from sensitivity to early exploration choices: different initial file sequences appear to trigger different architectural hypotheses, with some runs “locking in” to a more productive exploration path.
The belief externalization bottleneck.
The three Gemini models illustrate the belief externalization problem most clearly. Gemini 2.5 Flash opens 13 files (comparable to baselines) but recalled only 32% of IMPORTS edges, versus 55–58% for Config-Aware and Random. It appears to understand more than it reports: the model follows import chains when choosing exploration actions (evidence from action logs) but fails to transcribe observed imports into the cognitive map JSON. Gemini 3 Flash exhibits an extreme version: opening 12 files but reporting only 7 edges on average, with one codebase yielding just 4 edges total. This suggests that belief externalization—faithfully serializing internal understanding into a structured JSON schema—is itself a non-trivial capability. The gap between what an agent knows and what it writes down is a first-order confounder in any belief-probing benchmark.
Gemini 3.1 Pro: conservative but under-explored.
Gemini 3.1 Pro opens only 9 files on average (vs. 12–14 for other models), appearing to terminate exploration prematurely or allocate budget to LIST/SEARCH actions rather than OPEN. Despite high precision (0.894), its low file coverage directly limits recall (0.156). This suggests that the model’s exploration strategy—not its comprehension ability—is the primary bottleneck.
Precision-recall profiles differ qualitatively.
Baselines using AST parsing (BFS-Import, Random) achieve perfect precision—every reported edge is correct. Among LLMs, Claude Sonnet 4.6 also achieves near-perfect precision (0.991) while attaining the highest recall, suggesting that strong models can externalize beliefs both accurately and comprehensively. In contrast, Gemini 2.5 Flash has the lowest precision (0.568), generating 19 false CALLS_API edges and 8 false DATA_FLOWS_TO edges alongside its correct discoveries. Config-Aware shows a parallel pattern: perfect IMPORTS precision but only 33% REGISTRY_WIRES precision due to aggressive config-based inference.
Efficiency strongly favors LLM agents.
Under the Action AUC metric—which rewards early discovery—Claude Sonnet 4.6 dominates (0.377), nearly doubling Config-Aware (0.212). All five LLMs except Gemini 3 Flash (0.120) outperform BFS-Import (0.079) on AUC. This reflects the models’ ability to form useful architectural hypotheses from the first few files, while rule-based methods require more observations before deterministic parsing yields comparable coverage.
Invariant discovery is uniformly zero.
All methods score Invariant F1 = 0.0. LLM agents do report invariants (Claude averages 8.0, Gemini 2.5 Flash 6.0, GPT-5.3-Codex 5.7 per codebase), but none match the planted ground-truth constraints under structured form comparison. The Oracle baseline also scores zero: it provides dependency edges only, and invariant discovery requires natural-language comprehension of test assertions and documentation that no current method attempts.
Random’s competitiveness and budget effects.
@@
Random exploration (F1 = 0.538) performs nearly as well as Config-Aware (0.577). This near-parity reflects the current scale: budget on 27–30 modules allows opening roughly half the files, making raw coverage a strong proxy for import discovery. A budget sweep (Appendix C) confirms that at , Config-Aware’s advantage grows to 3 over Random (0.175 vs. 0.056). The meaningful discrimination at lies in understanding what you’ve seen: LLM agents discover qualitatively different edge types despite comparable file coverage.

### 05 Mar 2026 v3

@@ Abstract
AI code agents excel at isolated tasks yet struggle with multi-file software engineering requiring architectural understanding. We introduce Theory of Code Space (ToCS), a benchmark that evaluates whether agents can construct, maintain, and update coherent architectural beliefs during codebase exploration. Agents explore procedurally generated codebases under partial observability—opening files under a budget—and periodically externalize their belief state as structured JSON, producing a time-series of architectural understanding. Three findings emerge from experiments with four baselines and six frontier LLMs. First, the Active-Passive Gap is model-dependent: one model builds better maps through active exploration than from seeing all files at once, while another shows the opposite—revealing that active exploration is itself a non-trivial capability absent from some models. Second, retaining structured belief maps in context acts as self-scaffolding for some models but not others, showing that the mechanism is model-dependent. Third, belief state maintenance varies dramatically: a smaller model maintains perfectly stable beliefs across probes while its larger sibling suffers catastrophic belief collapse—forgetting previously-discovered components between probes. We release ToCS as open-source software.111Code: https://github.com/che-shr-cat/tocs
@@ 1 Introduction
Our contributions: (1) the ToCS benchmark framework for active architectural belief construction in code (§3); (2) a procedural codebase generator with four typed edge categories and planted invariants (§4); (3) pilot experiments with four baselines and six LLMs showing a wide capability spread and three surprising findings—a model-dependent Active-Passive Gap, model-dependent self-scaffolding effects, and dramatic belief state instability differences across models (§6–6.5); (4) an open benchmark released for community evaluation (§7).
@@ 3.1 Problem Formulation
This paper evaluates Construct; Revise and Exploit are implemented in the framework but reserved for future evaluation (§7.5).
@@ 3.2 Environment and Action Space
The agent operates under a budget (default ). Critically, SEARCH returns only locations, never content—ensuring that architectural understanding requires deliberate OPEN decisions. INSPECT provides a lightweight discovery channel: docstrings often describe module purpose and relationships (e.g., “wraps stage X”), allowing agents to gather architectural hints at the cost of one action without reading full file contents.
@@ 3.3 Cognitive Map Probing
The externalized cognitive map contains: (1) component beliefs with status (observed/inferred/unknown), purpose, exported symbols with typed signatures, and typed dependency edges with per-element confidence; (2) invariant beliefs with structured canonical form (type, src, dst, via, pattern) and evidence pointers; (3) uncertainty tracking—explicit list of unexplored regions. The resulting time-series of maps (one every steps) captures how understanding develops, not just the final state.
@@ 3.4 Evaluation Modes
We report APG decomposition results for GPT-5.3-Codex in §6.4.
@@ 4.2 Four Edge Types
calls_api (17%): Runtime function calls between modules. Requires reading function bodies; docstrings may provide hints (e.g., “delegates to module X”).
@@
data_flows_to (7%): Data dependency where one module’s output feeds another’s input. Requires understanding orchestration logic; docstrings on the orchestrator may describe the data flow.
These proportions emerge from the Pipeline template: static imports dominate, while semantic edges arise only at architectural boundaries. Crucially, roughly one-third of edges are invisible to import-following, creating a meaningful gap between syntactic analysis and semantic understanding.
@@ 4.3 Invariant Planting
Each codebase contains 15–16 planted constraints across five types: boundary (forbidden dependencies), dataflow (required processing chains), interface (access-only-through-ABC), invariant (naming/structural conventions), and purpose (design rationales). Every constraint has a structured canonical form with five fields for machine-comparable scoring: type (constraint category), src and dst (the modules involved), via (the intermediary or mechanism, e.g., an ABC interface), and pattern (the structural rule, e.g., “no direct import”). For example, a boundary constraint might be (boundary, mod_a, mod_c, base.StageBase, ‘‘must access only through ABC’’). Each constraint has at least one evidence source (test file, structural pattern, or documentation).
@@ 5 Metrics
Compare predicted edges from the cognitive map against ground truth. An edge matches if (source, target, type) all agree via exact string matching. Targets must be specific file paths: a directory-level edge (e.g., registry.py stages/) does not match file-level ground truth (e.g., registry.py stages/mod_a.py). Similarly, symbol-qualified targets (e.g., base.py::StageBase instead of base.py) do not match even when the file path portion is correct. This strict matching is deliberate—our ground truth is component-level (one file = one node)—but means that models reporting correct relationships at the wrong granularity or format receive no credit. We report precision, recall, and F1 separately.
@@
Match agent-discovered constraints against planted constraints via structured form comparison on (type, src, dst, via), not text similarity. Strict matching requires exact agreement on all four fields. Relaxed matching normalizes paths (directory prefix stripped), treats empty ground-truth fields as wildcards, and uses greedy 1-to-1 assignment—accommodating models that identify the correct constraint type and endpoints but use slightly different path formats. We report relaxed F1 in the main tables; strict scores are near-zero for all models due to compounding format mismatches.
@@ 6 Experiments
We evaluate four rule-based exploration strategies and six frontier LLM agents from three providers on three generated codebases (seeds 42, 123, 999) under identical conditions: budget , probe interval .
@@ 6.1 Methods
Gemini 2.5 Pro: Google’s reasoning-optimized pro model (gemini-2.5-pro). Larger sibling of 2.5 Flash.
@@
Gemini 3.1 Pro: Google’s most capable model (gemini-3.1-pro-preview). The flagship of the Gemini 3 family.
All models receive identical prompts: a system prompt describing the partial observability environment and available actions, per-turn action prompts showing remaining budget and opened files, and structured JSON probe prompts with schema, per-type field semantics, and worked examples for each invariant type (Appendix A). Relative to an earlier prompt version, the current probe includes explicit edge type decision rules and pairwise invariant decomposition instructions. Temperature is set to 0 except for GPT-5.3-Codex (reasoning model constraint). All probe responses were parsed successfully via the deterministic repair parser (zero format failures across all runs).
@@ 6.2 Results
| Oracle | baseline | 1.000 | 1.000 | 1.000 | — | — | — |
| Config-Aware | baseline | 0.577 .014 | 0.736 | 0.475 | 0.212 .007 | 0.000 | 13 |
| Random | baseline | 0.538 .028 | 1.000 | 0.368 | 0.142 .008 | 0.000 | 13 |
| BFS-Import | baseline | 0.293 .060 | 1.000 | 0.173 | 0.079 .003 | 0.000 | 13 |
| GPT-5.3-Codex | LLM | 0.676 .059 | 0.782 | 0.597 | 0.306 .018 | 0.739 .067 | 15 |
| Claude Sonnet 4.6 | LLM | 0.664 .026 | 0.983 | 0.502 | 0.350 .005 | 0.778 .025 | 13 |
| Gemini 2.5 Flash | LLM | 0.470 .100 | 0.642 | 0.373 | 0.259 .037 | 0.517 .154 | 14 |
| Gemini 3.1 Pro | LLM | 0.321 .128 | 0.764 | 0.212 | 0.140 .049 | 0.423 .209 | 11 |
| Gemini 2.5 Pro | LLM | 0.242 .034 | 1.000 | 0.138 | 0.205 .027 | 0.376 .034 | 12 |
| Gemini 3 Flash | LLM | 0.147 .012 | 0.897 | 0.080 | 0.133 .016 | 0.125 .000 | 12 |
@@
| GPT-5.3-Codex | 0.69 | 0.15 | 0.31 | 1.00 |
| Claude Sonnet 4.6 | 0.56 | 0.15 | 0.00 | 1.00 |
| Gemini 2.5 Flash | 0.47 | 0.18 | 0.00 | 0.37 |
| Gemini 3.1 Pro | 0.27 | 0.00 | 0.50 | 0.00 |
| Gemini 2.5 Pro | 0.12 | 0.08 | 0.00 | 0.53 |
| Gemini 3 Flash | 0.11 | 0.00 | 0.00 | 0.05 |
@@ 6.3 Analysis
Two LLM agents surpass all baselines.
GPT-5.3-Codex (F1 = 0.676) and Claude Sonnet 4.6 (F1 = 0.664) both exceed Config-Aware (0.577) by 9–10 points. GPT favors broad coverage (highest recall, 0.597); Claude favors targeted exploration (near-perfect precision, 0.983). Both discover all ground-truth REGISTRY_WIRES edges.
@@
LLM agents collectively discover all four edge types (Table 3), while baselines find at most two. The most distinctive capability is DATA_FLOWS_TO discovery—requiring multi-hop reasoning through the orchestrator—where Gemini 3.1 Pro achieves 50% recall and GPT-5.3-Codex 31%. Claude found zero DATA_FLOWS_TO edges despite reading the runner code, illustrating an exploration-comprehension tradeoff: understanding the orchestrator is insufficient without also opening the linked stage modules.
Improved prompts unlock invariant discovery.
With improved probe prompts, LLM agents achieve substantial invariant F1 under relaxed matching: Claude leads (0.778), followed by GPT (0.739) and Gemini 2.5 Flash (0.517). All models scored zero under the earlier prompt version, confirming this was primarily a prompt specification problem (§7.3).
Belief state instability.
Gemini 2.5 Pro and 3 Flash reveal catastrophic belief state instability despite opening comparable numbers of files (Figure 2). Gemini 3 Flash exhibits recency bias—each probe reports only recently-examined components. Gemini 2.5 Pro builds a reasonable map (peak F1 = 0.33 at step 9), then destroys it in a single probe. In contrast, Gemini 2.5 Flash—the smallest model—loses zero correct edges across all probes. This suggests belief state maintenance is not a function of model scale but may depend on training objectives.
Precision-recall decoupling.
Claude Sonnet 4.6 and Gemini 2.5 Pro both achieve near-perfect precision (0.98), but with dramatically different recall (0.502 vs. 0.138)—high precision alone does not indicate strong understanding. GPT trades precision (0.782) for the highest recall (0.597).
Budget effects.
@@
Random exploration (F1 = 0.538) performs nearly as well as Config-Aware (0.577) at , but at Config-Aware’s advantage grows to 3 (Appendix C). The meaningful discrimination at lies in understanding what you’ve seen: LLM agents discover qualitatively different edge types despite comparable file coverage.
6.4 Active-Passive Gap Decomposition
We evaluate the APG decomposition (§3) on two models—GPT-5.3-Codex and Gemini 2.5 Flash—across all three codebases. Table 4 reports Dependency F1 under each condition.
| Model | Condition | Dep F1 | Inv F1 | Description |
| GPT-5.3 | Passive-full | 0.457 .147 | 0.757 .028 | All files at once |
| Passive-replay | 0.665 .034 | 0.752 .036 | Same trace, no decisions | |
| Active | 0.676 .059 | 0.739 .067 | Agent chooses actions | |
| Passive-oracle | 0.737 .043 | 0.683 .091 | Oracle-selected files | |
| Gem. 2.5F | Passive-full | 0.696 .125 | 0.696 .120 | All files at once |
| Passive-replay | 0.561 .141 | 0.665 .037 | Same trace, no decisions | |
| Active | 0.469 .099 | 0.516 .155 | Agent chooses actions | |
| Passive-oracle | 0.641 .058 | 0.349 .079 | Oracle-selected files | |
| APG decomposition (Dep F1): | ||||
| 0.219 / 0.226 | GPT: active full; Gemini: full active | |||
| 0.060 / 0.172 | Oracle helps both; larger gap for Gemini | |||
| 0.011 / 0.092 | GPT: 0; Gemini: passive-replay active | |||
The APG direction is model-dependent.
In Theory of Space (Zhang et al., 2026), models consistently degrade in active mode (). In code, the direction depends on the model. GPT-5.3-Codex exhibits : active exploration outperforms receiving the entire codebase at once, consistent across all three codebases. The likely explanation is information overload: presenting 27–30 files simultaneously overwhelms the model’s ability to identify dependency relationships, while sequential exploration allows focused processing. Gemini 2.5 Flash shows the opposite pattern (): passive-full substantially outperforms active (Dep F1 0.696 vs. 0.469). This suggests that Gemini benefits from seeing all files at once rather than exploring incrementally—consistent with its strong “thinking” reasoning capabilities but weaker exploration strategy.
Selection cost is positive for both models.
Passive-oracle exceeds Active for both GPT (+0.060) and Gemini (+0.172), indicating suboptimal file selection in both cases. The larger gap for Gemini suggests its exploration strategy is more inefficient, opening less informative files relative to the oracle ordering.
Decision cost reveals a qualitative difference.
For GPT-5.3-Codex, Active Passive-replay (0.676 vs. 0.665): the cognitive overhead of making exploration decisions is negligible. For Gemini 2.5 Flash, the pattern reverses: Passive-replay (0.561) exceeds Active (0.469) by 9 points (). When given GPT’s observation trace without having to make decisions, Gemini builds better maps than when exploring on its own. This “negative decision cost” suggests that Gemini’s exploration decisions actively hurt its comprehension—perhaps by directing attention to less informative files or by the cognitive overhead of action selection competing with architectural reasoning.
Invariant discovery patterns.
For both models, passive-full yields the highest invariant F1 (GPT: 0.757, Gemini: 0.696). Passive-oracle scores lowest for both (GPT: 0.683, Gemini: 0.349). Seeing all files at once provides the global context needed for cross-cutting constraint discovery, while sequential presentation—even with optimal file ordering—limits the model’s ability to synthesize invariants spanning multiple modules. Gemini’s invariant scores under passive-oracle are notably low (0.349), suggesting it struggles to combine evidence presented incrementally.
6.5 Probe Condition Ablation
We evaluate the effect of probe conditions on both GPT-5.3-Codex and Gemini 2.5 Flash by comparing three tracking modes: probe_as_scratchpad (baseline—JSON map retained in context), no_probe (final map only, no intermediate probes), and probe_only (JSON collected but stripped from context after each probe). Table 5 reports results across 3 codebases.
| Model | Track | Dep F1 | Inv F1 |
| GPT-5.3-Codex | Probe-as-scratchpad | 0.676 .059 | 0.739 .067 |
| No-probe | 0.538 .127 | 0.570 .191 | |
| Probe-only | 0.464 .056 | 0.447 .028 | |
| Gemini 2.5 Flash | Probe-as-scratchpad | 0.469 .099 | 0.516 .155 |
| No-probe | 0.480 .161 | 0.273 .040 | |
| Probe-only | 0.401 .143 | 0.481 .222 | |
| Scratchpad effect (scratchpad no-probe): | |||
| GPT-5.3-Codex | 0.138 | 0.169 | |
| Gemini 2.5 Flash | 0.011 | 0.243 | |
Scratchpad provides a substantial boost for GPT but not for Gemini.
For GPT-5.3-Codex, retaining probe JSON in context improves Dep F1 by 13.8 points over the no-probe baseline. The model uses its own prior maps as external working memory, scaffolding subsequent exploration and maintaining belief coherence across probes. This quantitatively confirms the mechanism described in L6 (§7): the probe is not a passive measurement but an active intervention that shapes exploration.
For Gemini 2.5 Flash, the scratchpad effect on Dep F1 is negligible (0.011): no-probe (0.480) and scratchpad (0.469) perform comparably. However, scratchpad provides a large boost to invariant discovery (0.243), suggesting that Gemini uses the retained map primarily for cross-cutting reasoning rather than dependency tracking. This model-dependent pattern reveals that the scratchpad mechanism works through different channels for different models.
Probe-only is consistently the weakest condition.
For both models, probe-only yields the lowest Dep F1 (GPT: 0.464, Gemini: 0.401). Producing structured output consumes attention and context budget without the compensating benefit of self-reference. For GPT, probe-only is even below no-probe (0.074), confirming that the overhead of generating structured JSON without retention is actively harmful. For Gemini, the gap is smaller but in the same direction (0.079).
Implications.
For GPT, the ordering is clear: scratchpad no-probe probe-only, with a 14-point F1 gap between scratchpad and no-probe. For Gemini, the ordering is no-probe scratchpad probe-only, with the primary scratchpad benefit appearing in invariant discovery rather than dependency tracking. This model-dependent pattern complicates benchmark design: the “best” probe condition is not universal. We retain probe-as-scratchpad as the default because it enables the richest trajectory analysis and provides the strongest results for the highest-performing model, but report all three conditions for transparency.